# LoSiA: Efficient High-Rank Fine-Tuning via Subnet Localization and Optimization

### v3

@@ 3.3 Subnet Optimization and Intergration
During fine-tuning, the locations of core subnets may undergo dynamic shifts, as illustrated in Figure 3. Although a small subset of neurons is consistently selected, peripheral components exhibit significant temporal variability. Fine-tuning with a static subnet risks model underfitting and neuron over-specialization. To address the issue, we introduce an asynchronous and periodic subnet re-localization mechanism that adapts to the evolving network topology.
Figure 3: Core Subnet Distribution during Training. The subnet various across different training iterations.
Naive periodic re-localization strategies can induce training instability and loss spikes Lialin et al. (2023). Furthermore, the storage requirements for $\overline{I}(\cdot),\overline{U}(\cdot)$ in a synchronous period would lead to a scaling of GPU memory overhead. Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate. Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each layer $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$ with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$. The training timeline is divided into time slots of $T$ steps, such that for time slots $iT,(i+1)T),i=1,2,\ldots$, we:
@@
    Compute $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in the time slot $(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
@@
    Sequentially reselect $S_{l}$ by $s(\cdot)$ before step $t=(kL+l)T$, the end of time slots in 1.
This yields a reselection period of $\overline{T}=LT$ for each layer. The asynchronous design ensures that at any step, there is exactly one decoder layer calculating $\overline{I}(\cdot),\overline{U}(\cdot)$ and one rewarming, which greatly reduces additional GPU memory overheads for importance score calculation. We also implement a learning rate rewarm-up to further enhance training stability. Formally, the learning rate at step $t$ is:
@@
The condition $C$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}\$, where $T_{w}$ is the warmup duration. Figure 4 illustrates the timelines of importance calculation and the learning rate rewarming across layers, with re-localization sandwiched between them.
@@ 3.3.1 Faster Implementation (LoSiA-Pro)
Through subnet fine-tuning, LoSiA can further mitigate activation storage and backward latency. The gradient for the subnet $S$ can be factorized as:
@@
Noticing $\tilde{L}_{S}\in\mathbb{R^{np\times bs}},\tilde{R}_{S}\in\mathbb{R}^{bs\times mp}$, the input activation storage is reduced by a factor $p$, while the computational complexity of gradient calculation is reduced from $O(nmbs)$ to $O(nmbsp^{2})$. We named the method LoSiA-Pro, a refined equivalent implementation of LoSiA. It offers a 27.6% latency reduction compared to LoRA with Gradient Check-Pointing, while reducing 13.4GB GPU memory consumption compared to LoSiA training without Gradient Check-Pointing.
@@ 4 Experiments
We evaluate LoSiA across a broad range of model scales and datasets, conducting rigorous comparisons with common baselines. On both domain-specific and common-sense reasoning tasks, the method demonstrates robust performance with significantly reduced training overheads. The experiments highlight that LoSiA effectively promotes both training efficiency and task proficiency.
@@ 4.1 Experimental Setup
Models are trained on downstream tasks in the domains of mathematics, coding, and general capabilities. Specifically, training sets are sampled by 50,000 random entries from MetaMathQA, Magicoder, and Alpaca-GPT4, respectively. The GSM8K, MBPP, and MMLU benchmark are for testing. Additionally, we also compared LoSiA with baseline methods on eight common sense reasoning tasks. More details regarding the datasets can be found in the Appendix.
Table 1: Comparison of PEFT Methods Across Models on Domain-Specific Tasks. Accuracy is reported, alongside with memory consumption (GB), maximum per-task training time (h). The numbers in parentheses indicate the training time of LoSiA-Pro, which is a refined and computationally equivalent implementation of LoSiA.
Model Method Mem(GB) Time(h) GSM8K MBPP MMLU Avg.
@@
[…]
LoRA $36.1$ $15.0$ $35.7$ $41.1$ $26.0$ $36.6$ $34.9$ $31.2$ $34.25$
PiSSA $36.1$ $14.9$ $38.5$ $46.5$ $26.4$ $39.0$ $33.8$ $32.6$ $36.13$
DoRA $37.3$ $29.8$ $39.7$ $43.0$ $31.4$ $\mathbf{43.2}$ $36.2$ $37.1$ $38.43$
GaLore $37.5$ $14.2$ $39.3$ $44.7$ $\mathbf{31.6}$ $42.6$ $36.6$ $35.5$ $38.38$
LoSiA (-Pro) $36.9$ $\mathbf{10.3\ (9.4)}$ $\mathbf{42.8}$ $\mathbf{49.7}$ $30.7$ $43.0$ $\mathbf{37.5}$ $\mathbf{37.4}$ $\mathbf{40.18}$
[…]
LoRA $23.7$ $33.3$ $42.9$ $46.7$ $26.0$ $37.8$ $42.3$ $37.3$ $38.83$
PiSSA $23.7$ $33.1$ $43.5$ $46.2$ $26.8$ $36.6$ $42.7$ $38.5$ $39.05$
DoRA $24.2$ $68.3$ $\mathbf{45.0}$ $\mathbf{47.2}$ $26.0$ $34.4$ $44.1$ $36.7$ $38.90$
GaLore $23.7$ $39.6$ $42.2$ $45.3$ $28.0$ $39.0$ $43.1$ $41.2$ $39.80$
LoSiA (-Pro) $21.9$ $\mathbf{26.1\ (21.8)}$ $44.7$ $46.7$ $\mathbf{28.4}$ $\mathbf{39.4}$ $\mathbf{45.0}$ $\mathbf{41.5}$ $\mathbf{40.95}$
[…]
LoRA $36.9$ $56.0$ $58.6$ $\mathbf{56.4}$ $34.1$ $44.8$ $52.6$ $53.7$ $50.03$
PiSSA $36.9$ $55.5$ $53.4$ $55.2$ $34.5$ $44.8$ $52.0$ $48.8$ $48.11$
LoSiA (-Pro) $36.9$ $\mathbf{46.5\ (38.6)}$ $\mathbf{59.0}$ $54.0$ $\mathbf{34.9}$ $\mathbf{48.2}$ $\mathbf{53.1}$ $\mathbf{55.7}$ $\mathbf{50.82}$
@@
Additionally, both GaLore and LoSiA incorporate the output layer into the fine-tuning process. Dimension reduction factor $p_{0}$ is set to $\frac{1}{64}$ for Gemma 2B, $\frac{1}{8}$ for LLaMA-2 7B, and $1$ for LLaMA-2 13B in LoSiA. The PEFT modules are applied to all linear layers within the transformer. The training batch size is set to 4, the warm-up ratio is set to 0.1 and the model is trained by 3 epochs. For training stability, the backbone model are in precision of BF16 and low-rank modules are upcasted to FP32.  [^2] Trained with LLaMA-Factory Zheng et al. (2024). Upcasting to FP32 only costs an additional 0.6GB of memory and trains faster than BF16 in practice. All of the experiments are conducted on single NVIDIA A800 80GB GPU. Further details (including implementation details for common-sense reasoning tasks) can be found in Appendix A.3.
@@ 4.2 Main Results
Table 1 presents the performance of LoSiA compared to baseline methods across Gemma-2B, LLaMA2-7B, and LLaMA2-13B models. For GSM8K, we report 0-shot Chain-of-Thought (CoT) and 5-shot accuracy to reveal the model’s reasoning capability and few-shot prompting performance. For MBPP, we report the Pass@1 and Pass@10 metrics. For MMLU, we report both 5-shot generation and perplexity-based results. The metrics are intended to measure the quality of generation and knowledge proficiency, respectively. Table 2 shows results on common-sense reasoning tasks, extracting the option with minimum perplexity and reporting ACC metric following lm-evaluation-harness. The test setup provides a robust measure of intrinsic knowledge acquisition.
@@
LoSiA demonstrates superior knowledge retention, as evidenced by perplexity-based evaluations. It outperforms LoRA by 2.48% on commonsense reasoning tasks and maintains an average 1.93% improvement on MMLU (0-shot, PPL). Unlike low-rank methods, LoSiA’s sparse, high-rank fine-tuning approach enables localized knowledge retention while shifting likelihood toward correct answers.
@@
Figure 5 compares the training overheads of various PEFT methods. Contrast to baselines such as DoRA which incur significant additional FLOPs, LoSiA shows superior efficiency in both training time and memory usage. By eliminating extra matrix multiplication operations, LoSiA achieves faster training speeds. Its refined implement, LoSiA-Pro, further compresses activation storage by at least 22.8GB (w GC) and reduce the training time up to 34% (w/o GC) compared to LoRA by saving and computing on partial activations. A detailed training latency and GPU memory measurement is in Appendix A.4.
@@ 4.3 Ablation Study
This section assesses the functionality of sensitivity importance-aware localization, asynchronous mechanism, and re-warmups, alongside with robustness analysis of LoSiA. We present comprehensive ablation studies in Table 3 and training dynamics in Figure 6. Additional robustness tests for rank factor selection are provided in the Appendix.
Table 3: Ablation Study of LoSiA on GSM8K and MMLU, using LLaMA-2 7B as the backbone.
@@
Asynchronous mechanism yields more stable training
Variant SL refers to using a synchronous layer-wise localization mechanism. However, it causes loss fluctuation, destabilizes later training, and degenerate the model performance by 1.36% in average, while asynchronous updates produce more stable loss curves.
@@
Variant GL uses absolute gradients as the importance score. On MMLU, its performance remains comparable but is biased toward Humanities tasks (see Table 11), while its accuracy on GSM8K drops by 1.66%. Sensitivity-based scores, which aggregate multi-sample information, are more effective to capture general patterns in linear layers compared to biased gradients. However, gradient-based LoSiA exhibits promising results. In practice, the storage of $\overline{I}(\cdot),\overline{U}(\cdot)$ (about 1GB memory occupation on LLaMA-2 7B) can be eliminated using gradient-based importance if needed. Further discussion is provided in Appendix A.2.2.
@@
Table 4: Robustness of Time Slot $T$ Across a Series of Data Scales. Trained by MetaMathQA and evaluated by GSM8K on LLaMA-2 7B.
@@
Table 4 assesses the performance of LoSiA across different training data scales. LoSiA consistently outperforms LoRA, demonstrating stability and robustness. Furthermore, the optimal time slot $T$ is positively correlated with the size of training set, while LoSiA shows transcendent performance within a reasonable range of $T$.
@@ 4.4 Analysis
We analyze the core subnet selection frequency of neurons in Figure 7. The frequently selected neurons remain similar under different rank factor $p$, while smaller $p$ produces more concentrated distribution patterns. This indicates that LoSiA effectively identifies and optimizes critical neurons with limited training budgets, while simultaneously adjusting marginal parameters to further enhance generalization capability.
@@
To evaluate LoSiA’s efficacy in continual learning, we perform sequential fine-tuning on Hellaswag, PiQA, BoolQ, SiQA, and Winogrande datasets on LLaMA-2 7B. We employs Average Performance (AP) Chaudhry et al. (2018), Forward Transfer (FWT) Lopez-Paz and Ranzato (2017), and Backward Transfer (BWT) Ke and Liu (2023) to assess overall performance, knowledge transfer ability from previous tasks to current task, and level of forgetting, respectively. Details of the metrics and experiments are provided in A.3.4.

### v4

@@ 3.3 Subnet Optimization and Intergration
During fine-tuning, the locations of core subnets undergo dynamic shifts, as illustrated in Figure 3.
Figure 3: Core Subnet Distribution during Training. The chosen subnets various across training iterations.
Although a small subset of neurons is consistently selected, peripheral components exhibit significant temporal variability. Freezing a fixed mask therefore invites under-fitting and over-specialization of the lucky prophase winners. To address the issue, we introduce an asynchronous periodic subnet re-localization mechanism that adapts to the evolving network topology.
Naive periodic learning strategies can induce training instability and loss spikes Lialin et al. (2023). Furthermore, the storage requirements of $\overline{I}(\cdot),\overline{U}(\cdot)$ for every layer simultaneously would lead to a scaling of GPU memory overhead. Therefore, we propose asynchronous periodic localization coupled with rewarmings of learning rate techniques. Consider a model $f$ with $L$ decoder layers $\{D_{l}\}_{l=0}^{L-1}$, where each decoder $D_{l}$ contains $K$ linear layers $\{W_{l,k}\}_{k=1}^{K}$, with corresponding core subnets $\{S_{l,k}\}_{k=1}^{K}$. The training timeline is chopped into time slots of length $T$, such that for time slots $iT,(i+1)T),i=1,2,\ldots$, we:
@@
    Calculate $\overline{I}(\cdot),\overline{U}(\cdot)$ for layer $D_{l}$ in time slots $(kL+l-1)T,(kL+l)T),k\in\mathbb{N}$.
@@
    Sequentially reselect $S_{l}$ by $s(\cdot)$ before step $t=(kL+l)T$, the end of time slots.
Consequently, every core subnet is refreshed exactly once every $\overline{T}=LT$ steps, and, at any moment, only one layer is (i) accumulating importance statistics and (ii) rewarming learning rate. This greatly reduces the extra GPU memory footprint for importance score calculation.
The rewarming mechanism resets the learning rate to a short warm-up schedule to enhance training stability. Formally, the learning rate at step $t$ is:
@@
The condition $Cond$ is $t\in[(kL+l)T,(kL+l+1)T)\text{ and }t>T_{w}\$, where $T_{w}$ is the warmup duration. This means that rewarmings are triggered only after the initial warmup phase is finished. Figure 4 illustrates the timelines of importance calculation and the rewarming procedure across multiple layers. Importance scores are evaluated, immediately followed by learning rate rewarming, with re-localization sandwiched between them.
@@ 3.3.1 Faster Implementation (LoSiA-Pro)
Through subnet fine-tuning, LoSiA can further mitigate activation storage and backward latency. The gradient of subnet $S$ can be factorized as:
@@
Noticing $\tilde{L}_{S}\in\mathbb{R}^{np\times bs},\tilde{R}_{S}\in\mathbb{R}^{bs\times mp}$, the input activation storage is reduced by a factor $p$, while the computational complexity of the gradient calculation is reduced from $O(nmbs)$ to $O(nmbsp^{2})$. We named the method LoSiA-Pro, a refined equivalent implementation of LoSiA. It offers a 27.6% latency reduction compared to LoRA, while additionally reducing 13.4GB GPU memory consumption compared to LoSiA when training without Gradient Check-Pointing.
@@ 4 Experiments
We evaluate LoSiA across a wide range of model scales and datasets, conducting rigorous comparisons with common baselines. On both domain-specific and common-sense reasoning tasks, the method demonstrates robust performance with significantly reduced training overheads. The experiments highlight that LoSiA effectively promotes both training efficiency and task proficiency.
@@ 4.1 Experimental Setup
Models are trained on downstream tasks in the domains of mathematics, coding, and general capabilities. Specifically, training sets are sampled by 50,000 random entries from MetaMathQA, Magicoder, and Alpaca-GPT4, respectively. The GSM8K, MBPP, and MMLU benchmarks are for testing. Additionally, we also compared LoSiA with baseline methods on eight common sense reasoning tasks. More details regarding the datasets can be found in the Appendix.
Table 1: Comparison of PEFT Methods Across Models on Domain-Specific Tasks. Accuracy is reported, alongside with memory consumption (GB) and per-token training latency ($\mu s$ / token). The numbers in parentheses indicate the latency of LoSiA-Pro, which is a refined and computationally equivalent implementation of LoSiA.
Model Method Mem
Latency
($\mu s$ / token)
GSM8K MBPP MMLU Avg.
@@
Gemma 2B FFT $50.1$ $142.9$ $46.4$ $50.4$ $33.0$ $43.4$ $36.1$ $37.0$ $41.05$
LoRA $36.1$ $136.7$ $35.7$ $41.1$ $26.0$ $36.6$ $34.9$ $31.2$ $34.25$
PiSSA $36.1$ $136.9$ $38.5$ $46.5$ $26.4$ $39.0$ $33.8$ $32.6$ $36.13$
DoRA $37.3$ $296.6$ $39.7$ $43.0$ $31.4$ $\mathbf{43.2}$ $36.2$ $37.1$ $38.43$
GaLore $37.5$ $162.4$ $39.3$ $44.7$ $\mathbf{31.6}$ $42.6$ $36.6$ $35.5$ $38.38$
LoSiA (-Pro) $36.9$ $\mathbf{119.9\ (107.2)}$ $\mathbf{42.8}$ $\mathbf{49.7}$ $30.7$ $43.0$ $\mathbf{37.5}$ $\mathbf{37.4}$ $\mathbf{40.18}$
LLaMA 2-7B FFT $64.1$ 359.2 $46.6$ $46.9$ $29.9$ $40.2$ $45.2$ $42.5$ $41.88$
LoRA $23.7$ $338.0$ $42.9$ $46.7$ $26.0$ $37.8$ $42.3$ $37.3$ $38.83$
PiSSA $23.7$ $338.5$ $43.5$ $46.2$ $26.8$ $36.6$ $42.7$ $38.5$ $39.05$
DoRA $24.2$ $656.4$ $\mathbf{45.0}$ $\mathbf{47.2}$ $26.0$ $34.4$ $44.1$ $36.7$ $38.90$
GaLore $23.7$ $437.7$ $42.2$ $45.3$ $28.0$ $39.0$ $43.1$ $41.2$ $39.80$
LoSiA (-Pro) $21.9$ $\mathbf{290.4\ (244.8)}$ $44.7$ $46.7$ $\mathbf{28.4}$ $\mathbf{39.4}$ $\mathbf{45.0}$ $\mathbf{41.5}$ $\mathbf{40.95}$
LLaMA 2-13B FFT-8Bit $77.1$ $640.7$ $61.2$ $55.7$ $35.7$ $43.2$ $53.6$ $56.2$ $50.93$
LoRA $36.9$ $621.1$ $58.6$ $\mathbf{56.4}$ $34.1$ $44.8$ $52.6$ $53.7$ $50.03$
PiSSA $36.9$ $622.3$ $53.4$ $55.2$ $34.5$ $44.8$ $52.0$ $48.8$ $48.11$
LoSiA (-Pro) $36.9$ $\mathbf{548.5\ (453.4)}$ $\mathbf{59.0}$ $54.0$ $\mathbf{34.9}$ $\mathbf{48.2}$ $\mathbf{53.1}$ $\mathbf{55.7}$ $\mathbf{50.82}$
@@
Additionally, both GaLore and LoSiA incorporate the output layer into the fine-tuning process. Dimension reduction factor $p_{0}$ is set to $\frac{1}{64}$ for Gemma 2B, $\frac{1}{8}$ for LLaMA-2 7B, and $1$ for LLaMA-2 13B in LoSiA. The PEFT modules are applied to all linear layers within the transformer. The training batch size is set to $4$, the warm-up ratio is set to $0.1$ and the model is trained by $3$ epochs. For training stability, the backbone models are in precision of BF16 and low-rank modules are upcasted to FP32.  [^2] Trained with LLaMA-Factory Zheng et al. (2024). Upcasting to FP32 only costs an additional 0.60.6GB of memory. All of the experiments are conducted on single NVIDIA A800 80GB GPU. Further details of the experimental setting (including implementation details on common-sense reasoning tasks) can be found in Appendix A.3.
@@ 4.2 Main Results
Table 1 presents the overall performance of LoSiA compared to baseline methods across Gemma-2B, LLaMA2-7B, and LLaMA2-13B models. For GSM8K, we report 0-shot Chain-of-Thought (CoT) and 5-shot accuracy to reveal the model’s reasoning capability and few-shot prompting performance. For MBPP, we report the Pass@1 and Pass@10 metrics. For MMLU, we report both 5-shot generation and perplexity-based results. The metrics are intended to measure the quality of generation and knowledge proficiency, respectively. Table 2 shows the results on common-sense reasoning tasks, extracting the option with minimum perplexity, and reporting ACC metric following lm-evaluation-harness. The evaluation provides a robust measure of intrinsic knowledge acquisition.
@@
LoSiA demonstrates superior knowledge retention, as evidenced by perplexity-based evaluations. It outperforms LoRA by 2.48% on common-sense reasoning tasks and maintains an average 1.93% improvement on MMLU (0-shot, PPL). Unlike low-rank methods, LoSiA’s sparse, high-rank fine-tuning approach enables localized knowledge retention while shifting likelihood toward correct answers.
@@
Figure 5 compares the training overheads of various PEFT methods. Contrast to baselines such as DoRA that incur significant additional FLOPs, LoSiA shows superior efficiency in both training latency and memory usage. By eliminating extra matrix multiplication operations, LoSiA achieves faster training speeds. Its refined implement, LoSiA-Pro, further compresses activation storage by at least 22.8GB (w/o GC) and raise training throughput to 1.38x (w GC) compared to LoRA by saving and computing on partial activations. A detailed training latency and GPU memory measurement is in Appendix A.4.
@@ 4.3 Ablation Study
This section assesses the functionality of sensitivity importance-aware localization, asynchronous mechanism, re-warmings, and re-localizations, alongside with robustness analysis of LoSiA. We present comprehensive ablation studies in Table 3 and training dynamics in Figure 6. Additional robustness tests for rank factor selection are provided in the Appendix.
Table 3: Ablation Study of LoSiA on GSM8K and MMLU benchmark, using LLaMA-2 7B as backbone.
@@
w/o Re-localization
$42.76$ $43.81$ $43.29$
Asynchronous re-localization yields more stable training
When each layer fine-tunes with fixed core subnets as ReLO represents, it results in persistent under-fitting and marked drop in test accuracy, confirming that key parameters shift frequently during training, necessitating dynamic and periodic localization of the selected core sub-network to adapt in real time. Variant SL refers to using a synchronous layer-wise localization strategy. However, it causes loss fluctuation, destabilizes later training, and degenerates the model performance by 1.36% on average, while asynchronous updates produce more stable loss curves.
@@
Variant GL adopts absolute gradients as the importance score. On MMLU, its performance remains comparable to LoSiA but is biased towards humanities tasks (see Table 12), while its accuracy on GSM8K drops by 1.66%. Sensitivity-based scores, which aggregate multi-sample information, are more effective in capturing general patterns in linear layers compared to biased gradients. However, the gradient-based variant exhibits promising results. In practice, the storage of $\overline{I}(\cdot),\overline{U}(\cdot)$ (about 1GB of memory occupation on LLaMA-2 7B) can be eliminated using gradient-based importance if needed. Further discussion is provided in Appendix A.2.2.
@@
Table 4: Robustness of Time Slot $T$ Across a Series of Data Scales. Trained with MetaMathQA and evaluated by GSM8K on LLaMA-2 7B.
@@
Table 4 benchmarks LoSiA against LoRA as the training corpus grows. Across every scale, LoSiA consistently outperforms LoRA, demonstrating stability and robustness. Furthermore, the optimal time slot $T$ is positively correlated with the size of the training set, while LoSiA shows transcendent performance within a reasonable range of $T$.
@@ 4.4 Analysis
Figure 7 visualizes how often each neuron is chosen in core subnets during training. The frequently selected neurons remain similar under different rank factors $p$. The smaller $p$ merely sharpens the histogram: mass is pushed into the most important weights, so more radical compression does not discard salient parameters. LoSiA simultaneously adjusts marginal parameters to enhance generalization capability, yielding better generalization whenever the budget is tight.
@@
To evaluate LoSiA’s efficacy in continual learning, we perform sequential fine-tuning on Hellaswag, PiQA, BoolQ, SiQA, and Winogrande datasets on LLaMA-2 7B. We employ Average Performance (AP) Chaudhry et al. (2018), Forward Transfer (FWT) Lopez-Paz and Ranzato (2017), and Backward Transfer (BWT) Ke and Liu (2023) metrics to assess overall performance, knowledge transfer ability from previous tasks to current task, and level of forgetting, respectively. Details of the experiments are provided in Appendix A.3.4.
Suppose that the model learns sequentially on $N$ tasks. Let $P_{i,j}$ denote the accuracy on task $j$ after training on task $i$. Following Zhao et al. (2024b); Feng et al. (2025), we formulate the metrics (AP, FWT and BWT) as bellow:
Average Performance: The metric reflects overall task performance after continued learning, which is, $AP=\frac{1}{N}\sum_{i=1}^{N}P_{N,i}$
Backward Transfer: The metric evaluates the impact of learning later tasks on the model’s performance on an earlier task, that is, $\text{BWT}=\frac{1}{N-1}\sum_{i=1}^{N-1}(P_{N,i}-P_{i,i})$.